### 04 Mar 2026 v2

# OneRanker: Unified Generation and Ranking with One Model in Industrial Advertising Recommendation

Abstract.
The end-to-end generative paradigm is revolutionizing advertising recommendation systems, driving a shift from traditional cascaded architectures towards unified modeling. However, practical deployment faces three core challenges: the misalignment between interest objectives and business value, the target-agnostic limitation of generative processes, and the disconnection between generation and ranking stages. Existing solutions often fall into a dilemma where single-stage fusion induces optimization tension, while stage decoupling causes irreversible information loss. To address this, we propose OneRanker, achieving architectural-level deep integration of generation and ranking. First, we design a value-aware multi-task decoupling architecture. By leveraging task token sequences and causal mask, we separate interest coverage and value optimization spaces within shared representations, effectively alleviating target conflicts. Second, we construct a coarse-to-fine collaborative target awareness mechanism, utilizing Fake Item Tokens for implicit awareness during generation and a ranking decoder for explicit value alignment at the candidate level. Finally, we propose input-output dual-side consistency guarantees. Through Key/Value pass-through mechanisms and Distribution Consistency (DC) Constraint Loss, we achieve end-to-end collaborative optimization between generation and ranking. The full deployment on Tencent’s WeiXin channels advertising system has shown a significant improvement in key business metrics (GMV - Normal +1.34%), providing a new paradigm with industrial feasibility for generative advertising recommendations.
Recommender Systems, Generative Modeling, Advertising
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Learning to rankccs: Information systems Recommender systemsccs: Information systems Computational advertising
1. Introduction
Advertising and recommendation systems in large-scale internet applications are undergoing a fundamental transformation from traditional cascaded architectures to end-to-end generative paradigms. Driven by breakthroughs in large language models and sequence modeling techniques, this shift has fundamentally redefined how advertising platforms match users with relevant commercial content. As demonstrated by industrial systems such as the OneRec (Deng et al., 2025; Li et al., 2024a) series and GPR (Zhang et al., 2025), generative recommendation has emerged as a promising approach that formulates recommendation as an autoregressive generation task based on hierarchical semantic IDs, enabling direct optimization of final objectives while achieving high computational efficiency. However, as generative methods penetrate deeper into the ranking stage of advertising systems, three critical challenges have become central to the next-generation architecture design: unified modeling of interest objectives and business value within the generation process, precision degradation caused by target-agnostic generation, and the trade-off between consistency and computational efficiency.
Figure 1. Comparison between existing methods and ours.
First, there is a fundamental conflict between multi-token prediction (MTP) optimized based on user behavior (clicks/
conversions) and business value objectives (e.g., eCPM): existing solutions face a dilemma—single-stage fusion (e.g., directly injecting eCPM signals into MTP heads) causes interest coverage and value optimization to compromise each other within the shared representation space, inducing optimization tension, as shown in Figure 1(b); stage decoupling (generate-then-rank) prevents the generator from perceiving ranking objectives, leading to systematic filtering of high-value candidates at the generation stage, as shown in Figure 1(a). Second, the inherent target-agnostic limitation of generative models keeps user representations static during generation, unable to dynamically adapt to candidate item characteristics, severely constraining fine-grained user-item interaction capabilities. Third, the “generate-then-rank” architecture introduces a triple disconnection: representation inconsistency (heterogeneous representation spaces between generation and ranking), computational redundancy (repeated encoding of user representations), and error propagation (semantic drift in generation cannot be corrected during ranking), hindering the system from achieving global optimality. These challenges collectively constitute the key bottleneck preventing generative advertising recommendation from transitioning from “functional” to “excellent”.
To address these challenges, we propose OneRanker, whose core logic is shown in Figure 1(c). First, as shown in Figure 2, we introduce a value-aware multi-task decoupling architecture that achieves synergy rather than conflict between interest and value modeling. Specifically, we retain MTP’s original multi-interest heads to maintain coverage breadth while introducing a dedicated value-aware head for business value learning. Both components share underlying user representations for knowledge transfer but decouple output spaces through independent task tokens, effectively alleviating optimization tension between interest and value. Furthermore, we incorporate task ordering priors (e.g., impression $\rightarrow$ click $\rightarrow$ conversion $\rightarrow$ value) with causal mask, enabling high-level value tasks to dynamically absorb knowledge from low-level interest tasks, achieving progressive refinement from interest modeling to value optimization.
Secondly, to address the inherent target-agnostic limitation, we construct a coarse-to-fine collaborative target awareness mechanism. In Step 2 of Figure 2, we introduce Fake Item Tokens obtained by K-means clustering of the entire item space to enable coarse-grained implicit awareness, allowing the generation process to dynamically perceive item semantic distributions. Simultaneously, we design a dual-channel representation mechanism (task semantic channel + target aware channel) that naturally fuses semantic matching and target awareness through inner product operations during retrieval, explicitly enhancing target sensitivity in generation. In Step 3 of Figure 2, we introduce a ranking decoder that performs fine-grained target awareness through comprehensive cross-attention between candidate items and task tokens, enabling precise business value alignment.
Finally, to further enhance generation-ranking consistency and complete the unified framework, we propose dual-side consistency guarantees. On the input side, the ranking decoder constructs its Key/Value from both Step 1’s original multi-interest representations and Step 2’s refined representations, ensuring ranking decisions fully inherit generation process information. On the output side, we introduce a Distributional Consistency (DC) mechanism that propagates business value signals from Step 3’s ranker back to Step 2’s generator as soft labels. Through a joint loss function $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{MTP}}+\beta\mathcal{L}_{\text{rank}}+\gamma\mathcal{L}_{\text{consistency}}$, we achieve gradient-level collaborative optimization, enabling the generator to “anticipate” ranking preferences during training and fundamentally preventing objective drift between stages.
The main contributions of this work are summarized as follows:
  • We identify and alleviate the optimization tension between user interest modeling and business value optimization in generative advertising recommendation, proposing a value-aware multi-task decoupling architecture that achieves unified modeling of interest objectives and business value.
  • We construct a coarse-to-fine collaborative target awareness mechanism: coarse-grained awareness via Fake Item Tokens in the generation stage, and fine-grained awareness via the ranking decoder for precise business value alignment.
  • We design dual-side consistency guarantees through Key/Value fusion (input side) and joint loss optimization with DC (output side), enabling end-to-end collaborative optimization between generation and ranking.
  • We have fully deployed OneRanker on Tencent’s Weixin Channels advertising system. Large-scale online A/B tests demonstrate significant improvements on key business metrics (GMV-Normal +1.34%, Costs +0.72%), providing an industrially viable new paradigm for generative advertising recommendation.
2. Related Work
2.1. Value-Awareness in Generative Models
In recent years, the rise of generative models has shifted the recommender system paradigm from embedding-based matching to deep semantic understanding and generation. TIGER (Rajput et al., 2023) employs RQ-VAE (Lee et al., 2022) to produce quantized semantic identifiers, alleviating representation collapse in large-scale item spaces. GENRET (Sun et al., 2023) co-optimizes identifier generation with retrieval objectives, enabling semantic IDs to dynamically capture complex interaction patterns. OneRec (Deng et al., 2025) proposes a generative recommendation model based on MoE (Jacobs et al., 1991) and effectively improves sequence accuracy through DPO (Rafailov et al., 2023). Furthermore, OneRec-Think (Liu et al., 2025) introduces Chain-of-Thought (CoT) reasoning (Wei et al., 2022), significantly enhancing decision-making logic and interpretability in complex scenarios.
However, the unified modeling of user interest and business value in advertising ecosystems still faces endogenous challenges inherent to the generative paradigm. Unlike traditional discriminative models (Wen et al., 2020; Xi et al., 2021; Lin et al., 2019) that achieve objective alignment through loss weighting in parallel architectures, autoregressive generative frameworks suffer from two deep-seated contradictions: first, the long-chain gradient propagation in autoregression amplifies inter-task gradient conflicts; second, interest coverage and value optimization compete for distribution within a shared representation space. These limitations prompt us to transform multi-objective synergy from an ”external intervention” into an ”endogenous logic.”
2.2. Target-Awareness Limitations in Generative Models
Despite the rapid development of generative recommendation, existing generative methods still suffer from target-agnostic designs. In Transformer (Vaswani et al., 2017)-based models (Zhai et al., 2024; Kang and McAuley, 2018; Petrov and Macdonald, 2023), user history is often compressed into a single embedding, which lacks fine-grained interaction with candidate items. Consequently, the user representation remains constant during the generation process and cannot be dynamically adjusted based on the characteristics of candidate items. In advertising scenarios, where user intent is highly dynamic and influenced by multiple factors, this target-agnostic nature fundamentally limits the precision of identifying subtle user intentions from noisy behavioral signals.
To address this challenge, recent works have proposed several approaches: GRank (Sun et al., 2025) injects target signals into the generator during training; PROMISE (Guo et al., 2026) introduces a process reward model that computes matching scores with the target candidate space; and GRAM (Lee et al., 2025) ensures the decoding process is controlled in real-time by target attributes. However, these existing methods still suffer from coarse granularity or computational redundancy.
2.3. Joint Optimization of Generation and Ranking
Traditional ”Retrieval-Ranking” cascaded architectures have evolved under the generative paradigm into a two-stage ”Candidate Generation - Independent Ranking” workflow. In this process, the user’s historical behavior sequence serves as a context prefix to autoregressively generate candidate item tokens, which are subsequently processed by a separate ranking model for fine-grained scoring. However, this ”generation-then-ranking” paradigm often employs heterogeneous representation spaces, leading to semantic drift of user intent between stages and preventing the system from achieving global optimality.
In recent years, numerous studies have sought to advance the fusion of generation and ranking by integrating ranking signals directly into the generation process (Li et al., 2024b; Fan et al., 2025; Fu et al., 2026; Gao et al., 2025). While these proposed architectures imbue the generation process with ranking-level discriminative power, most remain at the ”post-processing” level of soft synergy and fail to fundamentally eliminate the heterogeneity of representation spaces. The OneRanker framework proposed in this paper achieves end-to-end hard synergy between generation and ranking by ensuring consistency across the input and output sides of both models, thereby eliminating error propagation between stages.
3. Methods
3.1. Overall Architecture
We propose OneRanker, an end-to-end generative advertising recommendation framework, as illustrated in Figure 2, which aims to systematically address three core challenges in generative advertising recommendation: (1) the objective misalignment between user interest modeling and business value optimization; (2) the inherent target awareness deficiency in the generation process; (3) the representation disconnection between generation and ranking stages. The framework consists of three logically progressive steps: Step 1 (Generation), we adopt the same tokenization method as GPR (Zhang et al., 2025) to convert user behavior sequences into heterogeneous token streams (including user tokens, context tokens, content tokens, and item tokens), and perform multi-interest path generation based on the HSTU Decoder-only architecture. The autoregressive Multi-Token Prediction (MTP) mechanism can generate multiple complete semantic ID paths in parallel within a single forward pass, which has become a standard operation in industrial generative recommendation systems. Step 2 (Multi-Task/Target-Aware), as the core innovation, decouples interest and value objectives through task tokens and introduces fake item tokens to achieve coarse-grained target awareness. Step 3 (Ranking), realizes business value alignment through fine-grained cross-attention mechanisms. The three steps are deeply coupled through Key/Value mechanisms, forming a collaborative optimization loop characterized by “value-guided generation, coarse-fine collaborative perception, and dual-side consistency guarantee”.
Figure 2. The overall framework of OneRanker, consists of three stages: (i) the Generation stage which utilizes a generative backbone network pre-trained with multiple tokens based on user behavior sequences using heterogeneous tokens; (ii) the Multi-Task/Target-Aware stage which decouples interest and value objectives via task tokens and enables coarse-grained target awareness via fake item tokens; (iii) the Ranking stage which achieves fine-grained value alignment through a unified decoder with input-output consistency guarantees;
3.2. Generation Enhancement: Multi-Task/Target-Aware Enhancement
This stage serves as an endogenous enhancement module within the generation pipeline, with its core objective being to enhance two critical capabilities of the MTP generation process: (1) multi-task decoupling capability—enabling different interest dimensions to perform specialized modeling for differentiated objectives, and separating interest modeling and business value optimization at the architectural level to alleviate their inherent optimization tension; (2) coarse-grained target awareness capability—mitigating the inherent “target-agnostic” limitation of generative models, enabling user representations to dynamically respond to candidate item characteristics. We achieve fine-grained calibration of multi-interest paths through the following four key designs, providing Step 3 with high-quality, target-aware input representations. This ensures that the final score alignment is built upon semantically rich and target-sensitive foundations, rather than “post-hoc correction” of low-quality generation results.
3.2.1. Value-Aware Multi-Task Decoupling Architecture
We achieve architectural level decoupling of interest modeling and business value optimization through a unified task token sequence, rather than mixing objectives within a single-head representation space. We construct a learnable task token sequence $\mathbf{T}=[\mathbf{t}_{i_{1}},\mathbf{t}_{i_{2}},\dots,\mathbf{t}_{i_{m}},\mathbf{t}_{v}]$, randomly initialized and optimized end-to-end during training. Specifically:
  • $\mathbf{t}_{i_{1}},\dots,\mathbf{t}_{i_{m}}$ are interest task tokens, focusing on modeling differentiated interest dimensions and capturing multi-interest distributions through independent prediction heads. Each token serves an independent interest objective or business task or jointly forms a sequential task chain (e.g., impression $\rightarrow$ click $\rightarrow$ conversion).
  • $\mathbf{t}_{v}$ is the value-aware task token, dedicated to learn business value (Final value).
All tokens share the underlying user representation but achieve task space decoupling through independent output heads, fundamentally alleviating the inherent tension between interest coverage and value optimization. The output at each task token position is routed to an independent prediction head, enabling explicit decoupling and parallel learning of multiple tasks. To model task dependencies, we propose the synergy of task ordering priors and causal mask: we introduce a task ordering prior—arranging tasks according to business sequence and cognitive difficulty (e.g., impression $\rightarrow$ click $\rightarrow$ conversion $\rightarrow$ business value)—while applying causal mask to ensure that subsequent tasks can access representations from preceding tasks, thereby enabling knowledge transfer and progressive refinement across tasks.
3.2.2. Fake Item Token: Coarse-Grained Target Awareness
To mitigate the inherent “target-agnostic” limitation of generative models (i.e. the lack of interaction between user representations and candidate items), we introduce fake item tokens $\mathbf{F}=[\mathbf{f}_{1},\dots,\mathbf{f}_{k}]$ as the core carrier for target awareness. These tokens are the $k$ cluster center vectors obtained by performing K-means clustering over the entire item space, representing core anchors in the item semantic space. In the input sequence, we concatenate the Fake Item Tokens after the Task Tokens to form a complete Query sequence $\mathbf{Q}=[\mathbf{T};\mathbf{F}]$; meanwhile, the output of Step 1 serves as the Key/Value for Cross-Attention. This design enables the model to dynamically perceive the item semantic distribution during the generation process, achieving coarse-grained implicit target awareness rather than relying solely on static user representations, thereby effectively alleviating the “target-agnostic” problem.
3.2.3. Heterogeneous Attention Decoder
We utilize the output of Step 1 as the Key/Value for the decoder, while the Query input comes from the constructed task tokens and fake item tokens $\mathbf{Q}=[\mathbf{T};\mathbf{F}]$. Notably, we introduce two key improvements over the standard Transformer Decoder:
  1. (1)
    Cross-Attention Prioritization: Traditional decoders first perform Self-Attention followed by Cross-Attention; we reverse this order. This design stems from a clear motivation: first aggregating and summarizing the user’s multi-interest representations (Key/Value from Step 1) through Cross-Attention to obtain sufficient user-item interaction information, and then refining the aggregated Query (Task + Fake Item) through Self-Attention for internal consolidation and inter-task collaboration. Experiments show that this order better aligns with the cognitive logic of “first understanding user intent, then refining task representations”, effectively enhancing representation quality.
  2. (2)
    Heterogeneous Mask Strategy: To ensure reasonable information flow and prevent information leakage, we design a refined mask mechanism:
    • Among Task Tokens: causal mask is applied to ensure temporal sequence modeling and dependency capture within the task sequence;
    • From Task Tokens to Fake Item Tokens: each Task Token is allowed to access all Fake Items, enabling implicit target-aware interaction;
    • From Fake Item Tokens to Task Tokens: each Fake Item is allowed to access all Task Tokens to fully capture multi-task semantics;
    • Among Fake Item Tokens: mutual invisibility is enforced to avoid interference between different cluster centers, ensuring each Fake Item independently performs “exploratory” matching with user intent.
3.2.4. Dual-Channel Representation Construction
The outputs of the heterogeneous attention decoder (Task Tokens and Fake Item Tokens) are fed into the MTP module to construct refined multi-interest representations. The core innovation lies in the dual-channel representation fusion mechanism:
  • Task Semantic Channel: Each Task Token $\mathbf{T}_{i}\in\mathbb{R}^{d_{\text{task}}}$ provides a semantic vector that encodes deep semantic representation for the task dimension.
  • Target Aware Channel: Each of the $k$ Fake Item Tokens $\mathbf{F}_{j}\in\mathbb{R}^{d_{\text{item}}}$ is concatenated with the current Task Token $\mathbf{T}_{i}$, then mapped through a dedicated MLP to produce a $k$-dimensional preference score vector $\mathbf{s}^{(i)}_{j}\in\mathbb{R}^{k}$ representing the user’s preference for Fake Item Token $\mathbf{F}_{j}$ in the current task $\mathbf{T}_{i}$. All k vectors are aggregated via sum pooling to form the target-aware representation:
    (1) $\mathbf{s}^{(i)}_{\text{target}}=\sum_{j=1}^{k}\mathbf{s}^{(i)}_{j}\in\mathbb{R}^{k}$
    which explicitly models the user’s preference distribution in the item semantic space under task $i$.
The final user representation for each MTP head $i$ is constructed by concatenating both channels:
(2) $\mathbf{e}_{\text{user}}^{(i)}=\text{Concat}\left(\mathbf{e}_{\text{task}}^{(i)},\,\mathbf{s}_{\text{target}}^{(i)}\right)\in\mathbb{R}^{d_{\text{task}}+k}$
This design achieves explicit target awareness: $\mathbf{s}^{(i)}_{\text{target}}$ serves as a “soft attention” over the item semantic distribution, providing fine-grained target-aware signals for subsequent retrieval.
To achieve symmetric enhancement, the item-side representation similarly incorporates a target-aware channel: given an item embedding $\mathbf{e}_{\text{item}}\in\mathbb{R}^{d_{\text{item}}}$, we compute its cosine similarities with all $k$ cluster centers $\{\mathbf{f}_{1},\dots,\mathbf{f}_{k}\}$ as $c_{i}=\text{cos\_sim}(\mathbf{e}_{\text{item}},\mathbf{f}_{i})$, and construct the symmetric enhanced item representation:
(3) $\mathbf{e}_{\text{item}}^{\text{enhanced}}=\text{Concat}\left(\mathbf{e}_{\text{item}},[c_{1},c_{2},\dots,c_{k}]\right)\in\mathbb{R}^{d_{\text{item}}+k}$
During retrieval, the relevance between user and item representations is computed via inner product:
(4) $\text{score}(user,item)=\mathbf{e}_{\text{user}}^{(i)}\mathbf{e}_{\text{item}}^{\text{enhanced}}=\mathbf{e}_{\text{task}}^{(i)}\mathbf{e}_{\text{item}}+\sum_{l=1}^{k}s^{(i)}_{l}\cdot c_{l}$
This inner product operation naturally fuses semantic matching with target-aware matching, endowing the retrieval process with explicit candidate item awareness and directly addressing the “target-agnostic” limitation of traditional generative models.
3.3. Unified Ranking
3.3.1. Fine-Grained Target Awareness
This stage is responsible for transforming the refined representations from Step 2 into the final ranking results. We design a dedicated ranking decoder (R-Decoder Block), whose core innovation lies in internalizing the ranking process as a natural extension of the generation pipeline, rather than an external post-processing module. Specifically:
Input Design: The Query of the R-Decoder consists of two parts: (1) a dedicated ranking task token $\mathbf{T}_{r}\in\mathbb{R}^{d_{\text{task}}}$, serving as a semantic anchor for ranking intent; (2) $n$ candidate item tokens $\mathbf{I}=[\mathbf{i}_{1},\dots,\mathbf{i}_{n}]$ from the multi-path generation results of Step 2 MTP, representing the candidate set to be ranked. The Key/Value fuses the original multi-interest representations from Step 1 with the refined representations from Step 2, ensuring that ranking decisions simultaneously leverage the user’s base interest distribution and the high-quality representations calibrated with target awareness.
Mask Strategy: Candidate item tokens $I$ are mutually invisible to each other (diagonal mask), avoiding interference among candidates and ensuring scoring independence for each candidate; All candidate item tokens $I$ can access the ranking task token ${T}_{r}$;
Architecture and Output: The R-Decoder adopts the same heterogeneous attention decoder structure as Step 2 (Cross-Attention prioritized followed by Self-Attention), but only use a single layer to complete ranking modeling. The output vector at each candidate item token position is directly mapped to a scalar score $s_{i}$ through a lightweight MLP. These scores $\mathbf{S}=[\mathbf{s}_{1},\dots,\mathbf{s}_{n}]$ naturally reside in a unified semantic space and can be directly used for final ranking without additional normalization or calibration.
3.3.2. Input-Side Consistency
Input-side consistency achieves seamless integration of the three stages through a Key/Value pass-through mechanism: the ranking decoder in Step 3 directly reuses the outputs of Step 1 and Step 2 as Key/Value, realizing efficient, and low-cost information flow connectivity.
This mechanism embodies dual semantic value:
  • Step 1 $\rightarrow$ Step 3: Transmits the user’s base interest distribution, providing breadth coverage guarantees for ranking;
  • Step 2 $\rightarrow$ Step 3: Transmits the hidden decision rationale of the generation process (i.e., the intermediate representations explaining “why this batch of candidates was generated”), enabling the ranker to fully comprehend the generation logic and achieve fine-grained reuse and calibration of the candidate set.
This design effectively alleviates the representation fragmentation problem inherent in traditional “external ranking” approaches—the ranker no longer faces “black-box” generation results, but can instead perceive the internal states of the generation process. Therefore, while maintaining the multi-benefit coverage advantage of MTP, the system has achieved further accuracy improvements.
3.3.3. Output-Side Consistency and Optimization Objectives
Output-side consistency aligns the optimization objectives of generation and ranking through a triadic joint loss function, enabling end-to-end collaborative optimization. The total loss is defined as:
(5) $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{MTP}}+\beta\mathcal{L}_{\text{rank}}+\gamma\mathcal{L}_{\text{DC}}$
where $\alpha,\beta,\gamma$ are the balance coefficients. $\mathcal{L}_{\text{MTP}}$ and $\mathcal{L}_{\text{rank}}$, respectively ensure the performance of the generation and ranking tasks, while $\mathcal{L}_{\text{DC}}$ is responsible for aligning the generation and ranking distributions.The components are defined as follows:
Generation Loss ($\mathcal{L}_{\text{MTP}}$): We employ a negative log-likelihood loss to optimize multi-interest path generation in Step 2. For each task head $j$ and semantic ID level $l$, the loss is formulated as:
(6) $\mathcal{L}_{\text{MTP}}=-\sum_{j=1}^{m}\sum_{l=1}^{L}\log P(\mathbf{e}_{\text{item}}^{\text{gt}}|\mathbf{e}_{\text{user}}^{(j)})$
where the conditional probability $P(\cdot)$ is constructed via softmax over the full semantic space $\mathcal{V}$ at that level:
(7) $\begin{split}P(\mathbf{e}_{\text{item}}^{\text{gt}}|\mathbf{e}_{\text{user}}^{(j)})=\frac{\exp(s_{j,\text{gt}})}{\sum_{k\in\mathcal{V}}\exp(s_{j,k})}\end{split}$
Here $s_{j,i}=\mathbf{e}_{\text{user}}^{(j)}\mathbf{e}_{\text{item}}^{(i)}$ represents the dot product score of user embedding and item embedding, the calculation details of which are shown in Equation 4. $\mathbf{e}_{\text{item}}^{\text{gt}}$ is the ground-truth semantic ID embedding for the next layer.
Specifically, while interest heads optimize for click/conversion matching, the value-aware head employs value-weighted sampling strategies to optimize the generation distribution towards high-value items, ensuring value propensity of generated candidates.
Ranking Loss ($\mathcal{L}_{\text{rank}}$): We adopt a pairwise BPR loss to optimize the relative ordering of candidates in Step 3:
(8) $\mathcal{L}_{\text{rank}}=-\mathbb{E}_{(i,j)\sim\mathcal{D}}\left[\mathbb{I}(y_{i}>y_{j})\cdot\log\sigma(s_{i}-s_{j})\right]$
where $(i,j)$ denotes an item pair, $y_{i},y_{j}$ are the real business value labels (eCPM) such that $y_{i}>y_{j}$ indicates item $i$ has higher business value than item $j$, and $s_{i}$ is the output of the Step 3 ranking decoder.
Distributional Consistency Loss ($\mathcal{L}_{\text{DC}}$): Although the value-aware head optimizes true value propensity, its signals are sparse and lack perception of the ranker’s calibrated scoring space. To bridge this, we treat the ranker as a teacher model and achieve decision boundary alignment by minimizing the KL divergence between generation and ranking preference distributions:
(9) $\mathcal{J}(\theta)=\mathbb{E}_{i\sim\pi_{\theta}}[s_{i}]-\lambda\cdot\text{KL}(\pi_{\theta}||\pi_{\text{ref}})$
where $\pi_{\theta}$ denotes the generator policy, $\pi_{\text{ref}}$ is the reference policy (original MTP model), $s_{i}$ represents the calibrated scores from the ranker, and $\lambda$ is the KL weight. To avoid instability from reinforcement learning sampling, we derive its supervised surrogate, the Distributional Consistency Loss:
(10) $\mathcal{L}_{\text{DC}}=-\mathbb{E}_{i\sim\mathcal{C}}\left[p_{i}^{\text{target}}\cdot\log\pi_{\theta}(i|\mathbf{u})\right]$
where $\mathcal{C}$ is the current candidate item set, $p_{i}^{\text{target}}=\text{softmax}(s_{i}/\tau)$ is the target probability distribution normalized from ranking scores, and $\tau$ is the temperature parameter. Unlike the local pairwise optimization of $\mathcal{L}_{\text{rank}}$, $\mathcal{L}_{\text{DC}}$ constrains global distributional consistency at the candidate-set level, establishing a dual-supervision mechanism of “ground-truth for propensity, model distribution for consistency.” This constraint establishes a backward gradient pathway from ranking to generation, enabling the generator to “anticipate” the ranker’s decision preferences and effectively correct value biases.
Through dual-side consistency guarantee, the OneRanker framework establishes a complete closed-loop optimization system: generation provides high-quality, target-aware candidates for ranking, while ranking supplies value-guided feedback signals for generation.
4. Experiments
To evaluate the performance and robustness of our proposed model, we conducted extensive experiments to address the following research questions (RQs):
  • RQ1: How does OneRanker perform in comparison with current strong generative recommendation baselines in the industry?
  • RQ2: How does the proposed key structural components contribute to the overall model performance?
  • RQ3: How effective are the other dedicated designs of the proposed architecture ?
  • RQ4: Does $\mathcal{L}_{\text{DC}}$ truly improve model consistency?
  • RQ5: What is the empirical performance and real-world gain of OneRanker in an online industrial environment?
4.1. Offline Experiments (RQ1)
4.1.1. Dataset Descriptions
Following the experimental setup in GPR (Zhang et al., 2025), we conduct our evaluation on a large-scale real-world dataset. This dataset comprises diverse user interactions across both advertising and organic content scenarios, ensuring robust representativeness and data scale. All historical behaviors are serialized in strict chronological order using a unified Heterogeneous Token Architecture (U/C/X/I), representing User, Content, Context, and Item tokens, respectively. For each target item, a specific set of candidates is sampled to construct training and evaluation instances for the business value learning task.
4.1.2. Baselines and Evaluation Metrics
We benchmark OneRanker against two strong generative recommendation baselines:
  • HSTU (Zhai et al., 2024): A trillion-parameter generative recommendation architecture developed by Meta, extensively deployed in large-scale industrial systems.
  • GPR (Zhang et al., 2025): An end-to-end generative recommendation framework proposed by Tencent; we utilize the MTP version as described in the original paper.
To comprehensively evaluate recommendation accuracy and ranking quality, we adopt Hit Ratio (HR@K) as our primary evaluation metrics.
4.1.3. Implementation Details
The embedding dimension for all token types is uniformly set to 128. In Step 1, the input U/C/X/I sequences are fixed to a maximum length of 2048 through truncation or padding to ensure input consistency, while the core G-Decoder module is configured with 4 Transformer layers. In Step 2, the number of interest-task tokens is set to 6 to align with the multi-task prediction heads. The value-aware task utilizes 2 specific task tokens whose outputs are aggregated via a bagging strategy. 32 Fake Item Tokens act as clustering centers to characterize the semantic distribution of items. The heterogeneous attention decoder consists of 2 layers. In Step 3, a single ranking-task token is used, and the R-Decoder depth is limited to 1 layer to achieve efficient inference and calibration.
4.1.4. Results
As illustrated in Table 1, we evaluate the performance of different architectures across various Hit Ratio (HR) metrics. Firstly, both HSTU and GPR serve as competitive baselines for capturing user interests. Notably, the GPR model demonstrates a clear advantage over HSTU across all measured intervals; for instance, it improves the HR@1 and HR@5 from 0.1741 and 0.4648 to 0.1824 and 0.4935, respectively. The proposed OneRanker model achieves a significant breakthrough, substantially outperforming both baselines by a wide margin. Specifically, OneRanker reaches an HR@1 of 0.2639, which represents a 44.7% improvement over the GPR model. Similar dominant gains are observed in HR@5 (0.6213) and HR@15 (0.8894). These results effectively validate that our integrated architecture, with its task-specific enhancements and value-aware ranking, can more precisely align model outputs with complex recommendation objectives, yielding superior performance across the board.
Table 1. Comparison of Experimental Results on Key Metrics
Model/Metric HR@1 HR@3 HR@5 HR@10 HR@15
HSTU 0.1741 0.3508 0.4648 0.6604 0.7953
GPR 0.1824 0.3703 0.4935 0.6957 0.8207
OneRanker 0.2639 0.4959 0.6213 0.7945 0.8894
4.2. Ablation Studies (RQ2&RQ3)
4.2.1. Ablation Studies on Key Structural Components (RQ2)
To quantify the contributions of various components, we designed different model variants to address the aforementioned research questions. The details of each variant are as follows:
  • OneRanker: The final version of the model, integrating all proposed features.
  • OneRanker w/o DC Loss: The OneRanker model with $\mathcal{L}_{\text{DC}}$ removed from Step 3.
  • OneRanker w/o S2 Info: The OneRanker model with all information from Step 2 (interest task tokens, value-aware tokens, and fake item tokens) removed from the Key/Value pairs, retaining only Step 1 information.
  • OneRanker with S3 Ranker Only: The Step 3 model with both the Distributional Consistency Loss and all Step 2 information removed. Key/Value pairs consist solely of Step 1 information.
  • OneRanker S2: A variant that removes Step 3 entirely, retaining all characteristics of Step 2.
  • OneRanker S2 w/o Target: The OneRanker S2 model with coarse-grained target-aware (Target) removed, retaining only the value-aware multi-task decoupled architecture.
  • OneRanker S2 w/o Target & MDA: The OneRanker S2 model with both Fake Item Token and the value-aware multi-task decoupled architecture (MDA) removed.
Table 2. Ablation Studies Results of Different Model Variants
Model/Method HR@1 HR@3 HR@5 HR@10 HR@15
OneRanker (Full) 0.2639 0.4959 0.6213 0.7945 0.8894
w/o DC Loss 0.2629 0.4925 0.6173 0.7905 0.8867
w/o S2 token injection 0.2610 0.4912 0.6161 0.7897 0.8868
with S3 ranker only 0.2601 0.4901 0.6157 0.7880 0.8852
OneRanker S2 0.2214 0.4257 0.5448 0.7291 0.8435
w/o Target 0.2080 0.4030 0.5203 0.7066 0.8276
w/o Target & MDA 0.2035 0.3922 0.5066 0.6949 0.8212
Experimental results indicate that each core module we designed brings significant improvements to model performance, and the ablation experiments further verify it, as is shown in Table 2. We primarily use HR@5 for comparison.
In Step 2, compared to the Base (OneRanker S2) HR@5 performance of 0.5448, the performance drops to 0.5203 (-4.5%) after removing Target information (w/o Target), which strongly proves the necessity of the Fake-Item-Token mechanism, prompting the model to evolve from static user representations to dynamic-aware item semantic distribution representations, thereby capturing users’ real-time intentions more accurately. Furthermore, removing both the Target information and the Multi-task Decoupled Architecture (w/o Target & MDA) results in an HR@5 decline to 0.5066 (-7.0%) compared to the OneRanker S2 base, underscoring the critical role of the decoupled architecture in mitigating optimization tension within the shared representation space, effectively preventing interference between business value modeling and user interest modeling.
In Step 3, using the OneRanker with S3 ranker only (HR@5: 0.6157) as the baseline, the model already demonstrates a leap-forward boost compared to S2-only models, proving that fine-grained target-aware mechanisms are fundamental to effective ranking. Building upon this, the introduction of S2 information injection (HR@5: 0.6161) ensures that the ranker reuses the generation logic to better calibrate its predictions within the ”generation context”. Furthermore, incorporating the Distributional Consistency ${L}_{\text{DC}}$ (HR@5: 0.6173) explicitly constrains the output distribution between stages, mitigating target shift through consistency regularization. Ultimately, the OneRanker (Full) model achieves the optimal HR@5 of 0.6213. This improvement stems from a collaborative effect: while S2 provides enriched feature injection to S3, the Distributional Consistency loss facilitates essential value-driven feedback from S3 back to S2. This closed-loop interaction between input consistency and output consistency allows the model to reach a higher ceiling of performance.
4.2.2. Ablation Studies on Dedicated Designs (RQ3)
We further investigate the effectiveness of several dedicated architectural design choices within Step 2. Specifically, we implement a Cross-Attention Prioritization (CA-Pri) Mechanism, which allows each task-specific token to sufficiently aggregate user information prior to refining task representations. Additionally, a Heterogeneous Mask Strategy is designed to regulate the information flow while preventing information leakage between disparate task tokens. As shown in Table 3, removing either the Cross-Attention Prioritization (CA-Pri) Mechanism or the heterogeneous masking (H-Mask) strategy from the Step 2 baseline leads to a significant performance degradation across all metrics. These results underscore the necessity of our refined attention mechanism and mask strategy in achieving superior recommendation accuracy.
Table 3. Ablation Studies on Dedicated Designs in Step 2
Model/Method HR@1 HR@3 HR@5 HR@10 HR@15
OneRanker S2 (Baseline) 0.2214 0.4257 0.5448 0.7291 0.8435
w/o CA-Pri 0.2117 0.4100 0.5277 0.7148 0.8337
w/o H-Mask 0.2149 0.4151 0.5335 0.7183 0.8368
4.3. Effectiveness Analysis of $\mathcal{L}_{\text{DC}}$ (RQ4)
To quantitatively evaluate the consistency between models with and without the Distributional Consistency Loss $\mathcal{L}_{\text{DC}}$ , we employ rank deviation boxplots and Top-K overlap curves across 30 candidate items, as is shown in Figure 3. Figure 3(a) characterizes the consistency by measuring the absolute rank differences of Step 2 and Step 3: the central box represents the interquartile range (IQR) defined by the first (Q1) and third (Q3) quartiles, while the internal horizontal line denotes the median and whiskers represent the minimum and maximum respectively. Figure 3(b) measures the alignment degree between the two steps when filtering the top $K$ items.
Experimental results demonstrate that incorporating the Distributional Consistency (DC) loss compresses the IQR (blue boxes), as evidenced by the markedly reduced box height compared to the baseline in Figure 3(a). This contraction reflects a high degree of decisional consistency of the two steps and enhanced stability. Correspondingly, the Top-K overlap curve in Figure 3(b) validates this distributional optimization: the optimized blue curve exhibits a higher starting point at $K=1$ and maintains a consistent lead until $K$ approaches the total candidate count. This confirms that the reduction in individual rank variance observed in the Figure 3(a) directly translates into stronger consistency effect, establishing a more robust multi-stage collaborative framework.
(a) Absolute Rank Difference
(b) Top-$K$ Item Overlap Ratio
Figure 3. Effectiveness of $\mathcal{L}_{\text{DC}}$: (a) Absolute rank difference of items between Step 2 and Step 3 at each rank for models with and without $\mathcal{L}_{\text{DC}}$; (b) Top-$K$ items overlap ratio between Step 2 and Step 3 with varying $K$ for the two models.
4.4. Online A/B Test (RQ5)
We evaluated and deployed OneRanker on Weixin Channels, a production advertising platform with hundreds of millions of active users and tens of millions of dynamic ads. The primary online metrics include GMV, GMV-Normal (ads optimized for clicks or conversions, constituting the majority of total GMV), and Costs, where GMV and GMV-Normal serve as the main Key Performance Indicators (KPIs) as they directly reflect business returns.
Our online baseline is a cascaded framework of generation + ranking: candidates generated by the generation model are fed into a separately trained ranking model to serve downstream applications. After rigorous A/B testing, OneRanker achieved stable and statistically significant improvements on the main KPIs compared to the previous version. The results are summarized in Table 4.
Table 4. Online A/B Test Results of OneRanker. Confidence intervals (CI) are calculated with 0.05 significance level. At 80% traffic phase, there is a phenomenon of traffic coverage.
Traffic Phase GMV GMV (CI) GMV-Normal GMV-Normal (CI) Costs Costs (CI)
5% +0.4067% [-0.8320%, 1.6454%] +1.3427% [0.1637%, 2.5216%] +0.7190% [0.4632%, 0.9747%]
20% +0.7796% [0.2849%, 1.2743%] +0.6446% [0.2139%, 1.0754%] +1.1462% [0.9958%, 1.2965%]
80% +0.0843% [-0.4531%, 0.6217%] +0.3500% [0.0188%, 0.6812%] +0.4493% [0.3261%, 0.5725%]
Based on these results, OneRanker was fully rolled out to 100% traffic and served as the production system for Weixin Channels advertising recommendation. The successful deployment validates OneRanker’s effectiveness in bridging the gap between generative modeling and business value optimization in large-scale advertising scenarios.
5. Conclusion
This paper presents the OneRanker, systematically addressing three core challenges in generative advertising recommendation: the misalignment between interest and value objectives, the target-agnostic limitation of generation processes, and the architectural disconnection between generation and ranking. By leveraging value-aware multi-task decoupling, coarse-to-fine collaborative target awareness, and input-output dual-side consistency guarantees, OneRanker achieves architectural-level deep integration and end-to-end collaborative optimization. Full-scale industrial deployment on Tencent’s Weixin Channels advertising system validates its effectiveness, yielding a +1.34% improvement in GMV-Normal. This work not only establishes a viable industrial paradigm but also completes the paradigm shift from ”stage decoupling” to ”architectural integration,” accumulating feasible experience for deploying generative technologies in complex business scenarios.

### 12 Mar 2026 v3

@@
 On the output side, we introduce a Distributional Consistency (DC) mechanism that propagates business value signals from Step 3’s ranker back to Step 2’s generator as soft labels.
-Through a joint loss function $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{MTP}}+\beta\mathcal{L}_{\text{rank}}+\gamma\mathcal{L}_{\text{consistency}}$, we achieve gradient-level collaborative optimization, enabling the generator to “anticipate” ranking preferences during training and fundamentally preventing objective drift between stages.
+Through a joint loss function $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{MTP}}+\beta\mathcal{L}_{\text{rank}}+\gamma\mathcal{L}_{\text{DC}}$, we achieve gradient-level collaborative optimization, enabling the generator to “anticipate” ranking preferences during training and fundamentally preventing objective drift between stages.
 The main contributions of this work are summarized as follows:
@@
   • GPR (Zhang et al., 2025): An end-to-end generative recommendation framework proposed by Tencent; we utilize the MTP version as described in the original paper.
-To comprehensively evaluate recommendation accuracy and ranking quality, we adopt Hit Ratio (HR@K) as our primary evaluation metrics.
+To comprehensively evaluate recommendation accuracy and ranking quality, we adopt Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K) as our primary evaluation metrics.
 4.1.3. Implementation Details
@@
 Firstly, both HSTU and GPR serve as competitive baselines for capturing user interests.
-Notably, the GPR model demonstrates a clear advantage over HSTU across all measured intervals; for instance, it improves the HR@1 and HR@5 from 0.1741 and 0.4648 to 0.1824 and 0.4935, respectively.
+Notably, the GPR model demonstrates a consistent advantage over HSTU across all measured intervals and metrics.
+For instance, GPR improves the HR@1 and NDCG@1 from 0.1741 and 0.6742 to 0.1824 and 0.6823, respectively.
 The proposed OneRanker model achieves a significant breakthrough, substantially outperforming both baselines by a wide margin.
-Specifically, OneRanker reaches an HR@1 of 0.2639, which represents a 44.7% improvement over the GPR model.
-Similar dominant gains are observed in HR@5 (0.6213) and HR@15 (0.8894).
-These results effectively validate that our integrated architecture, with its task-specific enhancements and value-aware ranking, can more precisely align model outputs with complex recommendation objectives, yielding superior performance across the board.
+Specifically, OneRanker reaches an HR@1 of 0.2639, which represents a 44.7% relative improvement over the GPR model.
+Similar dominant gains are observed in ranking quality metrics; for example, NDCG@5 increases from 0.6818 to 0.7904, and NDCG@15 increases from 0.7445 to 0.8206.
+These results effectively validate that our architecture, with its task-specific enhancements and value-aware ranking, can more precisely align model outputs with complex recommendation objectives, yielding superior performance across the board.
 Table 1. Comparison of Experimental Results on Key Metrics
-Model/Metric HR@1 HR@3 HR@5 HR@10 HR@15
-HSTU 0.1741 0.3508 0.4648 0.6604 0.7953
-GPR 0.1824 0.3703 0.4935 0.6957 0.8207
-OneRanker 0.2639 0.4959 0.6213 0.7945 0.8894
+Model/Method HR@1 HR@3 HR@5 HR@10 HR@15 NDCG@1 NDCG@3 NDCG@5 NDCG@10 NDCG@15
+HSTU 0.1741 0.3508 0.4648 0.6604 0.7953 0.6742 0.6712 0.6763 0.7025 0.7396
+GPR 0.1824 0.3703 0.4935 0.6957 0.8207 0.6823 0.6776 0.6818 0.7070 0.7445
+OneRanker 0.2639 0.4959 0.6213 0.7945 0.8894 0.8102 0.7954 0.7904 0.7970 0.8206
 4.2. Ablation Studies (RQ2&RQ3)
@@
 Table 2. Ablation Studies Results of Different Model Variants
-Model/Method HR@1 HR@3 HR@5 HR@10 HR@15
-OneRanker (Full) 0.2639 0.4959 0.6213 0.7945 0.8894
-w/o DC Loss 0.2629 0.4925 0.6173 0.7905 0.8867
-w/o S2 token injection 0.2610 0.4912 0.6161 0.7897 0.8868
-with S3 ranker only 0.2601 0.4901 0.6157 0.7880 0.8852
-OneRanker S2 0.2214 0.4257 0.5448 0.7291 0.8435
-w/o Target 0.2080 0.4030 0.5203 0.7066 0.8276
-w/o Target & MDA 0.2035 0.3922 0.5066 0.6949 0.8212
-Experimental results indicate that each core module we designed brings significant improvements to model performance, and the ablation experiments further verify it, as is shown in Table 2.
-We primarily use HR@5 for comparison.
-In Step 2, compared to the Base (OneRanker S2) HR@5 performance of 0.5448, the performance drops to 0.5203 (-4.5%) after removing Target information (w/o Target), which strongly proves the necessity of the Fake-Item-Token mechanism, prompting the model to evolve from static user representations to dynamic-aware item semantic distribution representations, thereby capturing users’ real-time intentions more accurately.
-Furthermore, removing both the Target information and the Multi-task Decoupled Architecture (w/o Target & MDA) results in an HR@5 decline to 0.5066 (-7.0%) compared to the OneRanker S2 base, underscoring the critical role of the decoupled architecture in mitigating optimization tension within the shared representation space, effectively preventing interference between business value modeling and user interest modeling.
-In Step 3, using the OneRanker with S3 ranker only (HR@5: 0.6157) as the baseline, the model already demonstrates a leap-forward boost compared to S2-only models, proving that fine-grained target-aware mechanisms are fundamental to effective ranking.
-Building upon this, the introduction of S2 information injection (HR@5: 0.6161) ensures that the ranker reuses the generation logic to better calibrate its predictions within the ”generation context”.
-Furthermore, incorporating the Distributional Consistency ${L}_{\text{DC}}$ (HR@5: 0.6173) explicitly constrains the output distribution between stages, mitigating target shift through consistency regularization.
-Ultimately, the OneRanker (Full) model achieves the optimal HR@5 of 0.6213.
+Model/Method HR@1 HR@3 HR@5 HR@10 HR@15 NDCG@1 NDCG@3 NDCG@5 NDCG@10 NDCG@15
+OneRanker (Full) 0.2639 0.4959 0.6213 0.7945 0.8894 0.8102 0.7954 0.7904 0.7970 0.8206
+w/o DC Loss 0.2629 0.4925 0.6173 0.7905 0.8867 0.8059 0.7912 0.7865 0.7936 0.8180
+w/o S2 token injection 0.2610 0.4912 0.6161 0.7897 0.8868 0.8047 0.7904 0.7858 0.7930 0.8172
+with S3 ranker only 0.2601 0.4901 0.6157 0.7880 0.8852 0.8037 0.7894 0.7849 0.7924 0.8168
+OneRanker S2 0.2214 0.4257 0.5448 0.7291 0.8435 0.7566 0.7460 0.7440 0.7579 0.7876
+w/o Target 0.2080 0.4030 0.5203 0.7066 0.8276 0.7375 0.7286 0.7285 0.7465 0.7786
+w/o Target & MDA 0.2035 0.3922 0.5066 0.6949 0.8212 0.7360 0.7273 0.7275 0.7453 0.7776
+Experimental results, as summarized in Table 2, demonstrate that each core module in our design contributes significantly to the overall performance.
+We primarily use HR@5 and NDCG@5 as the focal metrics for this ablation analysis to evaluate both hit accuracy and ranking quality.
+In Step 2 , compared to the Base (OneRanker S2) performance (HR@5: 0.5448, NDCG@5: 0.7440), removing the Target information (w/o Target) causes a notable decline, with HR@5 dropping to 0.5203 (-4.5%) and NDCG@5 decreasing to 0.7285 (-2.1%), which strongly proves the necessity of the Fake-Item-Token mechanism, prompting the model to evolve from static user representations to dynamic-aware item semantic distribution representations, thereby capturing users’ real-time intentions more accurately.
+Furthermore, the simultaneous removal of Target information and the Multi-task Decoupled Architecture (w/o Target & MDA) leads to a further decline to 0.5066 in HR@5 and 0.7275 in NDCG@5, underscoring the critical role of the decoupled architecture in mitigating optimization tension within the shared representation space, effectively preventing interference between business value modeling and user interest modeling.
+In Step 3, using the OneRanker with S3 ranker only (HR@5: 0.6157, NDCG@5: 0.7849) as the baseline, the model already demonstrates a leap-forward boost compared to S2-only models, proving that fine-grained target-aware mechanisms are fundamental to effective ranking.
+Building upon this, the introduction of S2 information injection (HR@5: 0.6161, NDCG@5: 0.7858) ensures that the ranker reuses the generation logic to better calibrate its predictions within the ”generation context”.
+Furthermore, incorporating the Distributional Consistency ${L}_{\text{DC}}$ (HR@5: 0.6173, NDCG@5: 0.7865) explicitly constrains the output distribution between stages, mitigating target shift through consistency regularization.
+Ultimately, the OneRanker (Full) model achieves the optimal HR@5 of 0.6213 and NDCG@5 of 0.7904.
 This improvement stems from a collaborative effect: while S2 provides enriched feature injection to S3, the Distributional Consistency loss facilitates essential value-driven feedback from S3 back to S2.